OpenStudy (anonymous):

If 3x + 4y=7 and 2x-3y=67 then which of the following is the value of xy?

OpenStudy (anonymous):

If xy is the product of x and y then, then multiply the x and y solution of the simultaneous equations in the problem statement. x y = 17*-11 = -187

OpenStudy (anonymous):

HOW DID YOU GET 17?=)

OpenStudy (anonymous):

The solution for x and y in solving the equations\[\{3x+4y\text{=}7, 2x-3y\text{=}67\}\]is 17 and -11 respectively. If the x and y values are plugged back into the equations\[\{3x+4y\text{=}7, 2x-3y\text{=}67\}\text{/.}\{\{x\to 17,y\to -11\}\}\]The Mathematica result is\[\{\{\text{True},\text{True}\}\} \]meaning that the LHS is equal to the RHs for each equation respectively.